### v1

# On the Conflict of Robustness and Learning in Collaborative Machine Learning

Abstract
Collaborative Machine Learning (CML) allows participants to jointly train a machine learning model while keeping their training data private. In scenarios where privacy is a strong requirement, such as health-related applications, safety is also a primary concern. This means that privacy-preserving CML processes must produce models that output correct and reliable decisions even in the presence of potentially untrusted participants. In response to this issue, researchers propose to use robust aggregators that rely on metrics which help filter out malicious contributions that could compromise the training process. In this work, we formalize the landscape of robust aggregators in the literature. Our formalization allows us to show that existing robust aggregators cannot fulfill their goal: either they use distance-based metrics that cannot accurately identify targeted malicious updates; or propose methods whose success is in direct conflict with the ability of CML participants to learn from others and therefore cannot eliminate the risk of manipulation without preventing learning.
1 Introduction
Collaborative Machine Learning (CML) allows users to jointly train a machine learning model. The main argument to favor CML instead of centralized learning techniques is privacy: CML enables training of models without the users’ data leaving their device. Users share learning proxies, e.g., gradients or model weights [fl], which are typically referred to as updates. CML architectures differ in the means to orchestrate updates amongst participants, with the main two approaches being server-aided (through Federated Learning [fl]), and peer-to-peer (through Decentralized Learning [dl]).
The privacy of these CML designs has been studied in depth [privacyDL]. In this paper, we focus on robustness –i.e., the reliability and correctness of the model output (e.g., prediction or classification result) – a property crucial in many fields where the privacy of CML is typically considered an advantage. For example, in health applications [health1, health2, health3, health4, health5, health6, health7, helath8, health9, health10], where incorrect diagnosis can threaten the life of patient; or in autonomous driving [fldriving, FLdeepdriving, feddrive, driving4, driving5, driving6, driving7] where wrong predictions threaten the safety of passengers, bystanders, and the environment. Even in fields where robustness is not essential for safety, robustness may be needed for economic reasons, such as avoiding manipulations that change content visibility in advertisement applications [ads], or change credit risk prediction [credit].
Lack of robustness can have two causes. First, from non-malicious failures, such as users having low-quality data or an unreliable network connection that prevents them from participating in the network. Second, it can be the result of attacks performed by malicious participants whose goal is to poison the final model and influence its behavior in some undesirable way.
In this paper, we evaluate the robustness of CML in the presence of an adversary. Following the literature, we assume this adversary can send arbitrary model updates to other participants during training e.g., [rfa, Section 3.2],  [krum, Section 2] [scclipping, Section 1][rofl, Section 2.1][trimmedmedian, Section 1][baffle, Section 3.C]. In this scenario, naive CML implementations, such as averaging all received updates [fl], do not provide robustness: the aggregation function processes its inputs in a predictable manner, and hence a malicious user can craft updates to arbitrarily influence the output of the aggregation [open].
Robust training techniques from the centralized setting such as examining data quality [clean, demon, fail, roni] or testing for adversarial examples [advexp], do not apply to CML. This is because in CML users do not share raw and they are not guaranteed to share a common distribution. To address this limitation, researchers have developed a number of CML-tailored sophisticated functions to improve robustness [rfa, trimmedmedian, tm2, scclipping, ogclipping, krum, bridge, bylan, order, cao, cert, romoa, autofed, mixing, byrdie, xie, ubar, localpoison, fltrust, svmdl, vote, rofl]. These functions enable participants to evaluate the model updates they receive, and reject those that are deemed detrimental to the final model, i.e., participants learn as if there were only honest users in the network. These functions have been referred to as robust aggregators [scclipping], byzantine resilient defenses [byrdie], scoring mechanism [krum], or even poisoning detectors [local].
While the efficiency of state-of-the-art robust aggregators against a selection of attacks has been demonstrated, so far there is yet no evaluation that discusses the effectiveness of these algorithms against an strategic adversary that would adapt to the defense nor of their impact on learning. This follows the typical trend in systems’ evaluation in which researchers iterate attacks and defenses to arrive to the state-of-the-art. This was the case in the search for robust algorithms in the centralized machine-learning literature [carlini], and this trend is repeating in the collaborative setting. There is a growing amount of attacks against state-of-the-art robust aggregators [local, scclipping, alie, imp, history, limits] and new robust aggregators protecting against state-of-the-art attacks [scclipping, local, tm2, ogclipping, history]. None of these works provide a formal proof of robustness guarantees against a strategic adversary. This lack of formalism is even more problematic in light of the recent findings which demonstrate how standard experimental setups can misrepresent robustness [pitfalls].
In this paper, we avoid this cat-and-mouse process by systematizing the study of CML robust aggregators, and directly characterizing the limits of collaborative robust learning in presence of malicious strategic adversaries. This characterization reveals a fundamental limit to robustness as robust aggregators currently define it: obtaining robustness in CML is in direct conflict with learning. In other words, allowing participants to learn always incur a risk of honest users being manipulated. Such risk might be acceptable in some applications, like in advertising; or inadmissible due to potential harms to users, like in health or self-driving applications.
Concretely, we make the following contributions:
  • We introduce a taxonomy for robustness in CML that systematizes the different kinds of robust aggregation techniques in the literature. We find that most robust aggregators belong to one of two categories: those that evaluate the robustness of updates based on their distance, e.g., L2 distance to a reference point; and those that evaluate the robustness of updates based on their behavior, e.g., error rate.
  • We show that distance-based robust aggregators that allow for a level of learning must also allow for adversarial manipulations of the same order of magnitude. This implies that distance is not an adequate proxy for evaluating the maliciousness of updates. We empirically demonstrate our findings in two realistic applications of CML: learning in healthcare and autonomous driving.
  • We demonstrate that when using behavior-based robust aggregators, the quality of the evaluation that users make is inversely proportional to their ability to learn from collaboration. Thus, the only users that can make good decisions and accurately reject malicious updates (and therefore learn in a robust manner) are those whose prior knowledge implies that they do not benefit much from participating in collaborative learning. Conversely, users who have the highest learning potential cannot produce accurate robustness evaluations, and remain vulnerable to manipulation despite their use of robust aggregation. We empirically validate our results on two example behaviour-based aggregators.
2 Systematization of Robustness Metrics
A CML system consists of $n$ users who hold private local data. User $i$ holds data $S_{i}$ formed by tuples of samples $x$ and labels $y$ sampled from the domain $\mathcal{D}=(\mathcal{X},\mathcal{Y})$. Users are connected according to a graph $\mathcal{G}$, in which nodes are users and edges represent connections between users. We call this graph topology. A common topology in CML is a star, used in Federated Learning [fl], in which all nodes communicate with each other via a central entity.
In CML, users initialize their models with common public parameters $m^{0}$ and iterate over the following three steps until a stop condition is met: (1) Local training: At the beginning of epoch $t$, users apply gradient descent on their local model parameters $m^{t}$ using their local data $S_{i}$, and compute an update $\theta_{i}^{t}$. (2) Communication: Users share their model update $\theta_{i}^{t}$ with their neighbors— or the server in the case of federated learning. (3) Aggregation: Users (or the server in federated learning) aggregate the updates, e.g., compute their average or their median. Users obtain an updated model $m^{t+1}$ using the result of their local aggregation. In federated learning, the server updates the global model and sends it to the users.
The primary goal of the collaboration is to enable users to obtain a model which performs better on the learning task than a model trained solely on their local data. In many cases, the application also requires this model to be robust, i.e., the participants must not be able to influence the model in any undesired way. While undesired is subjective and potentially relative to the application, in all cases the goal is to prevent manipulations that result on erroneous model outputs.
In CML, the main tool to ensure robustness of the final model is the use of robust aggregators during the training process. Robust aggregators aim to filter out detrimental updates before performing the aggregation, such that only honest users contribute to training. Robust aggregators rely on an evaluation metric to distinguish benign updates from malicious ones. We use $R(\theta;\mathcal{D})$ to refer to the evaluation of the update $\theta$ on the inputs from $\mathcal{D}$, using the metric $R$. The output of the evaluation is not necessarily restricted to $\{benign,malicious\}$ but can be a number that participants use in conjunction with a threshold value $\delta$, e.g., $R(\theta;\mathcal{D})<\delta$, to identify benign updates. Using a threshold makes the evaluation more flexible, e.g., to adapt the robustness evaluation to the learning task or application, or to the learning stage (e.g., use a stricter evaluation at the final steps of the training process).
A robust aggregator is a concrete instantiation of the metric $R$ and a means to evaluate the exchanged updates, i.e., each proposition of a metric $R$ is a proposition of a definition of robustness. The metric $R$ must have two properties. First, $R$ must capture the difference between benign and malicious updates such that the robust aggregator can effectively reject malicious updates: if an update does not influence the model in any way that leads to erroneous results, then $R(\theta;\mathcal{D})$ must be 0 (or below $\delta$, when using a threshold). Second, participants must be able to correctly compute $R$, even when they have access to a limited dataset and limited knowledge on the learning task.
2.1 Existing Approaches to Robustness
The metric $R$ should, ideally, be informative about the maximal error of the model $m^{t+1}$ resulting of the use of the update $\theta$ on inputs from the domain $\mathcal{D}$, i.e., $R(\theta,\mathcal{D})\doteq\underset{(x,y)\sim\mathcal{D}}{max}[m^{t+1}(x)-y]$. Yet, computing such metric in the collaborative setting is impossible because it requires computing the output of the model on all samples from $\mathcal{D}$. In addition to being a potentially unbounded computational task, individual users do not know the whole domain of inputs (by definition of the problem).
As a consequence, metrics $R(\theta,\mathcal{D})$ proposed in the literature resort to the evaluation of proxies that can be easily computed. These proxies include (but are not limited to): average accuracy of the update on a fixed test set [xie], invariance of prediction under input perturbation [invariance], smoothness of decision regions [advex] when the update is a model itself, or more generally the low norm of weights [l2] or L2-distance to the average of previously received updates [trilemma].
In general, the evaluation of an update $\theta$ is based on a comparison between this update and a reference that we denote as $\theta^{*}$. The reference can be a gradient or a model, typically of the same format as $\theta$. In the literature, we find that the reference $\theta^{*}$ is either fixed before training given prior knowledge of the participants, or computed dynamically per epoch given information available at the time of evaluation. In the latter case, the evaluator can use local knowledge (e.g., a local update), or knowledge obtained from the other participants’ in the network (e.g., the average of all updates, assuming that the majority of users is honest). The closer the update being evaluated ($\theta$) is to this reference ($\theta^{*}$), according to $R$, the more likely $\theta$ is to be benign. When $\theta=\theta^{*}$, $R(\theta,\mathcal{D})=0$.
In the CML literature, we find two classes of metrics used to assess ‘closeness’, which we formalize in the next two sections. Systematizing robust aggregators through classes of metrics allows us to reason about the properties of a class instead of individual robust aggregators. Such systematization also enables to evaluate the robustness of new aggregators without the need of an analysis from scratch.
2.1.1 Distance-based Evaluation
The most common class of robust aggregators are those that evaluate a characteristic of the update itself, e.g., the update weights  [rfa, trimmedmedian, tm2, scclipping, ogclipping, krum, bridge, bylan, order, cao, cert, romoa, autofed, mixing, byrdie, rofl, fltrust, ubar, sentinel]. Approaches in this class measure the closeness between the update $\theta$ and a reference $\theta^{*}$ using a distance metric $\Delta$ —i.e. a metric that has reflexivity, non-negativity, symmetry, and respects the triangle inequality. We refer to this class of robust aggregators as distance-based. We formalize them in Definition 2.1 and give examples of possible $\theta^{*}$ and $\Delta$ in Examples LABEL:ex:krumLABEL:ex:rofl, and LABEL:ex:scc.
Definition 2.1 (Distance-based robustness).
We say a robust aggregator is distance-based if its evaluation function can be expressed as:
$R(\theta,\mathcal{D})\doteq\Delta(\theta,\theta^{*})$
for a distance metric $\Delta$ and a reference update $\theta^{*}$.