### v2

# On the Conflict between Robustness and Learning in Collaborative Machine Learning

Abstract
Collaborative Machine Learning (CML) allows participants to jointly train a machine learning model while keeping their training data private. In many scenarios where CML is seen as the solution to privacy issues, such as health-related applications, safety is also a primary concern. To ensure that CML processes produce models that output correct and reliable decisions even in the presence of potentially untrusted participants, researchers propose to use robust aggregators to filter out malicious contributions that negatively influence the training process. In this work, we formalize the two prevalent forms of robust aggregators in the literature. We then show that neither can provide the intended protection: either they use distance-based metrics that cannot reliably identify malicious inputs to training; or use metrics based on the behavior of the loss function which create a conflict with the ability of CML participants to learn, i.e., they cannot eliminate the risk of compromise without preventing learning.
I Introduction
Collaborative Machine Learning (CML) allows users to jointly train a machine learning model. The main argument to favor CML instead of centralized learning techniques is privacy: CML enables training of models without the users’ data leaving their device. In CML, users share model updates that act as learning proxies, e.g., gradients or model weights [fl]. CML architectures differ in the means they orchestrate the sharing of updates amongst participants, with the main two approaches being server-aided (e.g., Federated Learning [fl]), and peer-to-peer (e.g., Decentralized Learning [dl]).
In many applications where the privacy of CML is typically considered an advantage, robustness is a critical property –i.e., the reliability and correctness of the model output (e.g., prediction or classification result). For example in health applications [health1, health2, health3, health4, health5, health6, health7, helath8, health9, health10], where incorrect diagnosis can threaten the life of patient; or in autonomous driving [fldriving, FLdeepdriving, feddrive, driving4, driving5, driving6, driving7] where wrong predictions threaten the safety of passengers, bystanders, and the environment. Even in fields where robustness is not essential for safety, it may be needed for economic reasons, such as avoiding manipulations that change content visibility in advertisement applications [ads] or change credit risk prediction [credit].
Lack of robustness may stem from either non-malicious failures, such as users having low-quality data or an unreliable network connection that prevents them from participating in the network; or from attacks performed by malicious participants whose goal is to influence the behaviour of the trained model(s) in some undesirable way. There is a number of CML-tailored sophisticated functions to improve robustness [rfa, trimmedmedian, tm2, scclipping, ogclipping, krum, bridge, bylan, order, cao, cert, romoa, driving5, mixing, byrdie, xie, ubar, localpoison, fltrust, svmdl, vote, rofl] – referred to as robust aggregators [scclipping], byzantine resilient defenses [byrdie], scoring mechanism [krum], or even poisoning detectors [local]. These functions enable participants to evaluate the model updates they receive, and reject those that are deemed detrimental to the final model, i.e., participants learn as if there were only honest users in the network.
The privacy of CML has been studied in depth [privacyDL], and there is a growing amount of attacks against state-of-the-art robust aggregators [local, scclipping, alie, imp, history, limits] and new robust aggregators protecting against state-of-the-art attacks [scclipping, local, tm2, ogclipping, history]. Yet, there is no formal study of robustness in collaborative scenarios. Existing work focuses solely in proving the efficiency of state-of-the-art robust aggregators against a selection of existing attacks, and does not discuss the effectiveness of these defenses against an strategic adversary that adapts to the detection algorithm. This follows the typical trend in systems’ evaluation in which researchers iterate attacks and defenses to arrive at the state-of-the-art, such as in the search for robust algorithms in the centralized machine-learning literature [carlini]. The lack of formalism and evaluation is even more problematic in light of the recent findings which demonstrate how standard experimental setups can misrepresent robustness [pitfalls].
In this paper, we provide the first formalization of the robustness problem in CML and use this characterization to be the first to provide insights on the limits of the most prevalent robustness approaches in CML: in both approaches, for honest participants to learn they must always incur a risk of being manipulated. Such risk might be acceptable in some applications, like in advertising; or inadmissible in others, like in health or self-driving applications due to potential harms to users. Concretely, we make the following contributions:
  • We systematize the different kinds of robust aggregation techniques in the CML literature. We find that most robust aggregators belong to one of two categories: those that evaluate the robustness of updates based on their distance, e.g., L2 distance to a reference point; and those that evaluate the robustness of updates based on their behavior, e.g., error rate.
  • We provide a game-based formalization of the learning and robustness goals in CML. This formalization enables to reason in a principled way about the adequacy of robust aggregators design approaches and the trade-offs between these two goals.
  • We demonstrate that distance-based robust aggregators that allow participants to learn from others must allow for adversarial manipulations of the same order of magnitude. This is because distance is not an adequate proxy for evaluating the maliciousness of updates.
  • We prove that when using behavior-based robust aggregators, the quality of the evaluation that users make is inversely proportional to their ability to learn from collaboration. Thus, the only users that can make good decisions and accurately reject malicious updates are those whose prior knowledge implies that they do not benefit much from participating in collaborative learning. Conversely, users who have the highest learning potential cannot produce accurate robustness evaluations, and remain vulnerable to manipulation despite their use of robust aggregation.
  • We empirically demonstrate the trade-off for both kinds of robust aggregators on state of the art aggregators, targeting fields where CML has been hailed as the solution to privacy.
II Robustness in Collaborative Learning
Collaborative learning. A Collaborative Machine Learning (CML) system consists of $n$ users who hold private local data. User $i$ holds data $S_{i}$ formed by tuples of samples $x$ and labels $y$ sampled from the domain $\mathcal{D}=(\mathcal{X},\mathcal{Y})$. Users are connected according to a graph $\mathcal{G}$, in which nodes are users and edges represent connections between users. We call this graph topology. A common topology in CML is a star, used in Federated Learning [fl], in which all nodes communicate with each other via a central entity.
In CML, users initialize their models with common public parameters $m^{0}$ and iterate over the following three steps until a stop condition is met: (1) Local training: At the beginning of epoch $t$, users apply gradient descent on their local model parameters $m^{t}$ using their local data $S_{i}$, and compute an update $\theta_{i}^{t}$. (2) Communication: Users share their model update $\theta_{i}^{t}$ with their neighbors— or the server in the case of federated learning. (3) Aggregation: Users (or the server in federated learning) aggregate the updates, e.g., compute their average or their median. Users obtain an updated model $m^{t+1}$ using the result of their local aggregation. In federated learning, the server updates the global model and sends it to the users.
Robustness Adversary: Goal and Capabilities. The primary goal of CML is to enable users to obtain a model which performs better on the learning task than a model trained solely on their local data. In many cases, the application also requires this model to be robust, i.e., the participants must not be able to influence the model in any undesired way. While undesired is subjective and potentially relative to the application, in all cases the goal is to prevent manipulations that result in erroneous model outputs.
We evaluate robustness when one, or more, of the participants are adversarial. We only consider as potential adversaries participants that send updates to other participants performing (robust) aggregation, e.g., to the federated server [fl] or to peers in decentralized learning [dl]. This excludes the scenario where a malicious federated server sends updates to users that accept it without further checks, as in that case, the server can already perform any adversarial manipulation by design.
The adversary aims to degrade the performance of the system, either in an untargeted way, e.g., reducing the performance of the model as in a denial of service attack; or in a targeted fashion, e.g., changing the output of some chosen inputs. Following the literature, we assume this adversary to be malicious, i.e., they can send arbitrary model updates to other participants during training e.g., [rfa, Section 3.2],  [krum, Section 2] [scclipping, Section 1][rofl, Section 2.1][trimmedmedian, Section 1][baffle, Section 3.C].
The precise capabilities of the adversary depend on the CML protocol and the threat model of the application [open]. At a minimum, the adversary, who participates in the training, has knowledge of the training public parameters (model architecture, learning rate) and of the updates it receives from its neighbors. Other capabilities such as the ability to participate in all epochs can vary depending on the CML system.
III Systematization of Robust Aggregators
In CML, the main tool to ensure robustness of the final model is the use of robust aggregators during the training process. Robust aggregators aim to filter out detrimental updates before aggregation, such that only honest users contribute to training. Robust aggregators rely on an evaluation metric to distinguish benign updates from malicious ones. We use $R(\theta;\mathcal{D})$ to refer to the evaluation of the update $\theta$ on the inputs from $\mathcal{D}$, using the metric $R$. The output of the evaluation is used in conjunction with a threshold value $\delta$ to identify benign updates, e.g., $R(\theta;\mathcal{D})<\delta$.
A robust aggregator is a concrete instantiation of the metric $R$ and a means to evaluate the exchanged updates, i.e., each proposition of a metric $R$ is a proposition of a definition of robustness. The metric $R$ must have two properties. First, $R$ must capture the difference between benign and malicious updates such that the robust aggregator can effectively reject malicious updates: if an update does not influence the model in any way that leads to erroneous results, then $R(\theta;\mathcal{D})$ must be below $\delta$. Second, participants must be able to correctly compute $R$, even when they have access to a limited dataset and limited knowledge on the learning task.
Robustness metrics. The metric $R$ should, ideally, be informative about the maximal error of the model $m^{t+1}$ resulting of the use of the evaluated update $\theta$ on the current model $m^{t}$ on inputs from the domain $\mathcal{D}$, i.e., $R(\theta,\mathcal{D})\doteq\underset{(x,y)\sim\mathcal{D}}{max}[m^{t+1}(x)-y]$. Yet, computing such metric in the collaborative setting is impossible because it requires computing the output of the model on all samples from $\mathcal{D}$. In addition to being a potentially unbounded computational task, individual users do not know the whole domain of inputs (by definition of the problem).
As a consequence, metrics $R(\theta,\mathcal{D})$ proposed in the literature resort to the evaluation of proxies that can be easily computed. These proxies include (but are not limited to): average accuracy of the update on a fixed test set [xie], invariance of prediction under input perturbation [invariance], smoothness of decision regions [advex] when the update is a model itself, or more generally the low norm of weights [l2] or L2-distance to the average of previously received updates [trilemma].
In general, the evaluation of an update $\theta$ is based on a comparison between this update and a reference that we denote as $\theta^{*}$. The reference can be a gradient or a model, typically of the same format as $\theta$. In the literature, we find that the reference $\theta^{*}$ is either fixed before training given prior knowledge of the participants, or computed dynamically per epoch given information available at the time of evaluation. In the latter case, the evaluator can use local knowledge (e.g., a local update), or knowledge obtained from the other participants’ in the network (e.g., the average of all updates, assuming that the majority of users is honest). The closer the update being evaluated ($\theta$) is to this reference ($\theta^{*}$), according to $R$, the more likely $\theta$ is to be benign. When $\theta=\theta^{*}$, $R(\theta,\mathcal{D})=0$.
Robustness metrics in the literature. We find two prevalent classes of metrics in the CML literature to assess ‘closeness’, which we formalize in the next two sections. Systematizing robust aggregators through classes of metrics allows us to reason about the properties of a class instead of individual robust aggregators. It also ensures that our results apply to new aggregators without the need of an analysis from scratch. We discuss classes of metrics not covered by our analysis in Section LABEL:sec:outofscope.
III-A Distance-based Evaluation
The most common class of robust aggregators are those that evaluate a characteristic of the update itself, e.g., the update weights  [rfa, trimmedmedian, tm2, scclipping, ogclipping, krum, bridge, bylan, order, cao, cert, romoa, driving5, mixing, byrdie, rofl, fltrust, ubar, sentinel]. Approaches in this class measure the closeness between the update $\theta$ and a reference $\theta^{*}$ using a distance metric $\Delta$ —i.e. a metric that has reflexivity, non-negativity, symmetry, and respects the triangle inequality. We refer to this class of robust aggregators as distance-based. We formalize them in Definition III.1 and give examples of possible $\theta^{*}$ and $\Delta$ in Section III-A1.
Definition III.1 (Distance-based robustness).
We say a robust aggregator is distance-based if its evaluation function can be expressed as:
$R(\theta,\mathcal{D})\doteq\Delta(\theta,\theta^{*})$
for a distance metric $\Delta$ and a reference update $\theta^{*}$.
III-A1 Examples of distance-based robust aggregators
Multi-KRUM [krum]. Multi-KRUM is a robust Federated Learning aggregator function run by the server. It uses the KRUM score as the robustness evaluation function $R$. Multi-KRUM is an example of an aggregator whose reference point $\theta^{*}$ dynamically computed at each epoch using network knowledge. The KRUM score of an update $\theta$ captures the cumulative L2-distance between this update and the users’ updates. The server computes $R(\theta,\mathcal{D})=score(\theta)=\sum_{\theta_{j}\in CU(\theta)}||\theta-\theta_{j}||_{2},$ where $CU(\theta)$ are the $n-f-2$ closest updates (CU) to $\theta$ from all received updates. The KRUM score can be expressed as $\Delta(\theta,\theta^{*})$ with $\Delta$ the L2-distance, and $\theta^{*}$ a function of users’ updates such that $\sum_{\theta_{j}\in CU(\theta)}||\theta-\theta_{j}||_{2}=||\theta-\theta^{*}||_{2}$. The Multi-KRUM robust aggregator selects the $n-f$ updates with lowest score, and outputs their average as result of the aggregation.
RoFL [rofl]. ROFL is another robust Federated Learning aggregator function run at the server. It defines the evaluation function $R$ to be the norm of the update. ROFL is an example of an aggregator that uses a fixed reference point $\theta^{*}$ decided a-priori. In RoFL, users construct cryptographic proofs that the norm of their update is below a threshold selected by the server. Encryption aside, the RoFL aggregation function can be instantiated with $\Delta$ the L2-distance, and $\theta^{*}$ the all 0-model, i.e., $R(\theta,\mathcal{D})=||\theta||_{2}$.
Self-Centered Clipping [scclipping]. Self Centered Clipping (SCC) is a peer-to-peer-specific robust aggregation scheme. It uses the L2-distance as evaluation function $R$, and is an example of an aggregator whose reference point is dynamically computed each epoch based on local knowledge. In SCC, users evaluate an update $\theta$ by computing the L2-distance between this update and their own local update $\theta_{i}$: $R(\theta,\mathcal{D})=||\theta-\theta_{i}||_{2}.$ This function can be expressed as $\Delta(\theta,\theta^{*})$ with $\Delta$ the L2-distance, and $\theta^{*}$ the user’s own update $\theta_{i}$. SCC differs from other evaluators in that it does not directly reject the updates for which $R(\theta,\mathcal{D})>\delta$, but clips them into an update $\theta^{\prime}=\texttt{min}(1,\delta/||\theta-\theta_{i}||)\cdot(\theta-\theta_{i})+\theta_{i}$ that satisfies $R(\theta^{\prime},\mathcal{D})\leq\delta$, and discards the rest of the update. The output of the aggregation is the average of the accepted (clipped) updates.
III-B Behavior-based Evaluation
A second common class of robust aggregators are those that evaluate the behavior of the model after integrating the received update (or the behavior of the update itself when directly evaluated on the model) on selected inputs [xie, localpoison, svmdl, vote, basil, baffle, bygars, ubar, sentinel]. These evaluators compare this behavior to a pre-determined expected behavior modeled by a function $\pi$, e.g., the expected loss of the model. The reference $\theta^{*}$ determines the expected behavior. $\theta^{*}$ itself does not necessarily need to be known. Sometimes only its behavior $\pi(\theta^{*},\mathcal{D})$ matters for the evaluation purpose. We refer to this category of robust aggregators as behavior-based. We formalize them in Definition LABEL:def:behavior and give examples of choice of $\theta^{*}$ and $\pi$ in Section LABEL:sec:examples-behavior.